OpenStudy (anonymous):

Let f(x) = 3x - 10. Find the value of f(-3). a. 21 b. -1 c. -19

OpenStudy (mertsj):

f(-3) = 3(-3) -10 f(-3) = -9-10 = -19

OpenStudy (anonymous):

f(x)=3x-10 f(-3)=3(-3)-10 =-9-10 =-19 the answer is c

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

You are welcome :)